OpenStudy (anonymous):

m=2/3, (7,5) find the point slope equation of the line. what is the equation of the line?

OpenStudy (kropot72):

The general formula for the point slope equation of a line is: \[y-y _{1}=m(x-x _{1})\] \[m=\frac{2}{3}\] \[y _{1}=5\] \[x _{1}=7\] Just plug these values into the general equation to find the point slope equation of the line.